20 p. #3 Find the solution to the differential equation ry2 – xy? x + xy4 satisfying the initial condition y(1) = -2
Answer #1

-2 * e moyennyt y?-ayz . .40 30032 og athya ( a -12 (1+y4) = xy?(9-12 City 47 ou du = 71717-yo hiyo ty by using Xariable seguy (1)=-2 * FILI+C 4*** 3 2 Ź - 5+1=4 1-3 1-3=( Ic=] Then © becomes Isregnized sobiliary cune / +43 = 22-27 -2 Cor) +3+y4=293.